OpenStudy (anonymous):

How would I solve this? The average value of v(t) e^t 2cos t = over the domain [1, 4] is approximately: F. −1.149 × 106 G. −2.840 X 105 H. 2.516 × 105 I. 1.688 J. 0.562

OpenStudy (anonymous):

so I go \[1/(4-1)\int\limits_{1}^{4} e ^{t} 2cost \]?

OpenStudy (anonymous):

That's right, not sure how to integrate that though.

OpenStudy (anonymous):

@lalaly could you show me how to integrate this?

OpenStudy (lalaly):

u = e^t du = e^tdt dv = costdt v = sint (e^t)(sint) - ∫ sint e^t dt u = e^t du = e^t dt dv = sint v = -cost dt (e^t)(sint) - ∫ sint e^tdt (e^t)(sin t) - [ (e^t)(-cos t) - ∫ -cost e^t dt ] (e^t)(sin t) + (e^t)(cos t) - ∫ cost e^t dt so ∫ cos(t e^t dt = (e^t)(sin t) + (e^t)(cos t) - ∫ cos(x) e^t dt ∫ cos(t e^t dt = (e^t)(sin t) + (e^t)(cos t) - ∫ cost e^t dt 2 ∫ cost e^t dt = (e^t)(sin t) + (e^t)(cost ∫ cost e^t dt = [ (e^t)(sint) + (e^t)(cos t) ] / 2

OpenStudy (lalaly):

just plug in the limits from 1 to 4

OpenStudy (lalaly):

yeah and multiply the integral by 1/3 lol

OpenStudy (anonymous):

what's dv?

OpenStudy (lalaly):

i integrated by parts ,,, so u=e^t dv= costdt du=e^tdt v=sint